# LittleBit: Ultra Low-Bit Quantization via Latent Factorization

### v3

@@ 1 Introduction
Large Language Models (LLMs) based on the Transformer architecture [attention2017] have revolutionized natural language processing. However, their immense scale, reaching hundreds of billions or trillions of parameters  [efficient2021, scaling2023], leads to prohibitive computational and memory costs, particularly GPU VRAM, hindering their widespread deployment on consumer or edge devices [edge2024, instinfer2024, dovetail2024].
Model quantization, which reduces numerical precision, is a key technique to address these bottlenecks [billm2024]. While post-training quantization (PTQ) methods like GPTQ [gptq2023] and AWQ [awq2023] compress to $\sim$4 bits, further compression (e.g., to 1-bit) typically requires quantization-aware training (QAT) [qat2018, qakd2019] to maintain performance. Indeed, QAT has demonstrated the ability to enable substantial 1-bit compression while preserving model fidelity [bitnet2023, onebit2024, fbillm2024].
However, even 1-bit models (e.g., $\sim$15.4GB for 70B parameters [billm2024]) can exceed the capacity of highly resource-constrained devices [reviewEdgeLLM2024], motivating the exploration of sub-1-bit quantization. Although prior work [stbllm2024] has ventured into this area, maintaining performance at extremely low effective bits (e.g., 0.55 bits per weight (BPW)) while ensuring computational efficiency without restrictive hardware dependencies remains a critical challenge [stbllm2024]. Consequently, it is crucial to pioneer new techniques that can achieve such profound model compression, aiming for bit rates around 0.1 BPW, while maintaining the integrity of model performance.
@@
Our approach stems from two observations. First, LLM weight matrices often exhibit significant low-rank structure [galore2024, dobi2025]. This suggests that factorization methods like Singular Value Decomposition (SVD) [svd1970] could offer a more stable compression pathway than pruning [deepcompression2016], especially under extreme compression regimes. For instance, at compression ratios exceeding 50%, SVD-based approaches [svdllmv2_2025] consistently outperform pruning methods like Wanda [wanda2023] (see Section˜A.5 for details). Second, binarization inherently causes information loss [survey2024]. Recent high-performing 1-bit methods [onebit2024, binarymos2024, arbllm2025] demonstrate that sophisticated scaling over multiple dimensions (e.g., row and column), is crucial for mitigating this loss and stabilizing performance [arbllm2025].
@@
Extensive experiments show LittleBit’s superiority over STBLLM [stbllm2024], the leading sub-1-bit technique, on LLMs from 1.3B to 32B parameters. Notably, LittleBit achieves competitive performance on Llama2-13B at an unprecedented 0.1 BPW (Fig.˜1), surpassing STBLLM at 0.55 BPW. Furthermore, the 32B LittleBit model at just 0.3 BPW maintains strong performance, significantly outperforming STBLLM and setting a new state-of-the-art for sub-1-bit quantization.
@@
  • We provide extensive validation showing that LittleBit significantly outperforms STBLLM [stbllm2024], the prior leading sub-1-bit quantization technique, on various LLM scales.
@@ 2.1 Binarization and Quantization
Network binarization targets extreme compression and acceleration by constraining weights or activations to $\pm 1$, enabling efficient bitwise operations [binarized2016, towards2018]. Although early works like BinaryConnect [binaryconnect2015] and BNNs [binarized2016] showed feasibility, they faced significant accuracy degradation [binarysurvey2020]. This degradation was partially mitigated by introducing scaling factors [xnor2016] and employing the Straight-Through Estimator (STE) [ste2013] to handle the non-differentiable quantization function during training [custom2024].
However, directly applying these early techniques to sensitive large language models (LLMs) typically results in unacceptable performance loss [pbllm2023]. To address this challenge, LLM-specific quantization strategies have been developed, mainly categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [survey2021]. PTQ adapts pre-trained models, achieving competitive results around 4-bit precision with methods such as GPTQ [gptq2023] and AWQ [awq2023], but generally struggles to maintain performance below 2-bit precision [pbllm2023, billm2024]. The sub-1-bit regime is particularly challenging for PTQ methods due to severe information loss [predicting2013]. Recent work such as STBLLM [stbllm2024] shows that incorporating structured binarized compression can push PTQ into the sub-1-bit regime, yet a non-negligible accuracy gap remains.
In contrast, QAT integrates the quantization process directly into the training loop. This enables the model to learn and adapt to the low-precision format during training, generally yielding superior performance compared to PTQ at very low bit-widths [llmqat2023]. Successful QAT approaches for low-bit scenarios often employ sophisticated scaling mechanisms alongside the adaptive training itself, as evidenced by several recent methods [onebit2024, binarymos2024]. However, despite these advancements, consistently achieving high model fidelity when quantizing to less than 1.0 BPW remains a considerable challenge. This persisting difficulty underscores the need for novel architectures and training strategies specifically designed for such extreme compression regimes.
@@ 2.2 Low-Rank Approximation in Quantization
Beyond quantization, the inherent parameter redundancy within deep neural networks [predicting2013], particularly evident in LLMs [shortgpt2024, whatmatters2024], can be effectively exploited using low-rank approximation methods. Techniques such as Singular Value Decomposition (SVD) [svd1970] allow large weight matrices to be represented as products of smaller factors, offering an alternative compression strategy that can be more resilient than pruning, especially under high compression regimes [deepcompression2016]. Recognizing their potential for complementary benefits, researchers have explored combining low-rank factorization to reduce parameter count and quantization to lower parameter precision [caldera2024]. Initial studies applied SVD in various roles within the LLM context: some used activation statistics to guide decomposition [svdllmv2_2025], others integrated SVD principles into initialization or parameter-efficient updates [adalora2023, pissa2024, onebit2024], applied transformations prior to factorization [asvd2023], or even employed low-rank structures to model quantization errors [lqer2024].
@@
A more integrated approach involves incorporating low-rank concepts directly within the QAT methods, as demonstrated by DL-QAT [dlqat2024], which jointly learns a representation amenable to both low-rank approximation and low-precision quantization, albeit limited to 3-bit quantization. This direction of optimizing structural properties such as rank and numerical precision during training represents a promising path towards more effective LLM compression, with potential for further exploration of even lower precision quantization.
@@ 3.1 LittleBit Architecture
The LittleBit architecture, depicted in Fig.˜2, redesigns linear layers for extreme compression (e.g., 0.1 BPW) by synergistically combining low-rank factorization and binarization. LittleBit first leverages the observed low-rank structure in LLM weights [svdllmv1_2025] to approximate the weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ as:
@@
To compensate for the significant information loss incurred by binarization [survey2024], LittleBit incorporates learnable FP16 scales. Beyond standard row ($\mathbf{h}$) and column ($\mathbf{g}$) scaling [pbllm2023, onebit2024, arbllm2025], it introduces an additional latent scale ($\bm{\ell}\in\mathbb{R}^{r}$). This latent scale is designed to address the factorization structure by learning the relative importance of each of the $r$ latent dimensions, which correspond to the columns of the binarized factors $\mathbf{U}_{\mathrm{sign}}$ and $\mathbf{V}_{\mathrm{sign}}$:
@@ 4.1 Settings
We evaluate LittleBit on diverse LLM families including Llama [llama2023], Llama2 [llama2_2023], Llama3 [llama3_2024], OPT [opt2022], Phi-4 [phi4_2024], and QwQ [qwq32b2025], spanning sizes from 1.3B to 32B parameters. The primary performance metric is perplexity (PPL), evaluated on the WikiText-2 [wiki2_2016] validation dataset. Additional results on the C4 [C42020] and PTB [ptb1994] datasets are provided in Section˜A.3. We also assess zero-shot accuracy on common reasoning benchmarks: WinoGrande [winogrande2021], OpenBookQA (OBQA) [OpenBookQA2018], HellaSwag [hellaswag2019], BoolQ [boolq2019], ARC-Easy (ARC-e), ARC-Challenge (ARC-c) [arc2018], and PIQA [piqa2020].
@@
The LittleBit model parameters, initialized using the Dual-SVID method (Section˜3.2), are optimized using QAT with knowledge distillation (KD) [distilling2015, llmqat2023, ternarybert2020]. The original pre-trained full-precision model serves as the teacher ($\mathcal{T}$) for the LittleBit student model ($\mathcal{S}$). The QAT objective combines the standard output Kullback-Leibler (KL) divergence loss $\mathcal{L}_{\mathrm{out}}$ and an intermediate layer Mean Squared Error (MSE) loss $\mathcal{L}_{\mathrm{inter}}$ to match hidden representations, weighted by an empirically determined $\lambda=10$:
@@
Following the setup in [binarymos2024], the training data combine WikiText-2 and selected partitions from C4. We use 2048 token sequence length, 5 epochs, Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$[adam2014], and cosine LR decay with 2% warm-up (see details in  LABEL:appendix:hyper_parameters). For models employing Grouped-Query Attention (GQA) [gqa2023] (e.g., Llama3, Phi-4, QwQ), the latent ranks $r$ for key(K) and value (V) projection layers were specifically adjusted (see Section˜A.2.3) to maintain performance under extreme quantization. For backpropagation, SmoothSign (forward: $\mathrm{sign}(x)$; backward gradient: $\tanh(kx)$, $k=100$) is used, which generally performed better than Straight-Through Estimator (STE) [ste2013] (see Section˜A.2.2). While the concept of a proxy gradient is not new, our contribution lies in the specific application and empirical validation of using the derivative of $\tanh(100x)$ as a stable gradient for QAT in the ultra-low-bit regime, which our ablation study confirms is effective (Section˜A.2.2).
@@
We benchmark LittleBit across a range of effective bits from approximately 1.0 down to 0.1 bits per weight (BPW). In the challenging sub-1-bit regime, we compare against STBLLM [stbllm2024], a post-training quantization (PTQ) method employing N:M sparsity. For the $\sim$1.0 BPW setting, we compare against the 1-bit PTQ method BiLLM [billm2024] and quantization-aware training (QAT) approaches (OneBit [onebit2024], BinaryMoS [binarymos2024]). Results from standard low-bit methods like GPTQ [gptq2023] and OmniQuant [omniquant2023] are included for broader context.
Table 1: Perplexity (PPL) comparison on WikiText-2 across various LLMs and quantization methods. Lower PPL indicates better performance. BPW denotes effective bits per weight. Methods marked with use quantization-aware training (QAT). STBLLM [stbllm2024] utilizes N:M sparsity (ratio in parentheses). LittleBit demonstrates strong performance, particularly in the sub-1-bit regime.
@@
Table 2: Zero-shot accuracy (%) comparison on common sense reasoning benchmarks. Compares Full Precision (FP16) models against STBLLM [stbllm2024] and LittleBit at 0.55 and 0.3 BPW.
@@ 4.2 Main Results
Table˜1 shows LittleBit’s perplexity (PPL) results against baselines. LittleBit consistently performs robustly, especially in the sub-1-bit regime. Compared to STBLLM [stbllm2024], LittleBit achieves markedly superior PPL at 0.8, 0.7, and 0.55 BPW. For instance, on Llama2-7B, LittleBit at 0.55 BPW achieves a PPL of 10.47, a substantial improvement over STBLLM’s 30.67. Similar significant gains are observed across other models and BPWs; for Llama2-13B at 0.8 BPW, LittleBit records a PPL of 8.52 compared to STBLLM’s 11.90.
@@
A key strength of LittleBit is its exceptional stability and performance in the extreme sub-0.5 BPW range, specifically at 0.3 BPW and 0.1 BPW. In this territory, methods like STBLLM exhibit severe performance degradation (e.g., Llama2-7B PPL of 1.8e3 at 0.3 BPW). In contrast, LittleBit maintains strong PPL scores, such as 12.00 for Llama2-7B at 0.3 BPW and 15.92 at an unprecedented 0.1 BPW. When configured for an effective bits around 1.0 BPW by adjusting the latent rank $r$, LittleBit delivers perplexity comparable to specialized 1-bit QAT methods like OneBit [onebit2024] (e.g., Llama2-7B: LittleBit 9.08 vs. OneBit 8.36) and PTQ methods such as BiLLM [billm2024]. While a method like BinaryMoS [binarymos2024] may exhibit slightly lower PPL in some 1-bit scenarios (e.g., Llama2-7B: BinaryMoS 7.74), this can be attributed to its use of more complex mechanisms like dynamic scaling, which represent a different set of architectural and computational trade-offs. Analyzing the trade-off by adjusting latent rank to tune effective bits from 1.0 down to 0.1 BPW reveals that while performance degradation is minimal down to  0.55 BPW, a quantization cliff appears between 0.3-0.1 BPW, positioning the 0.3-0.55 BPW range as an optimal sweet spot for balancing compression and accuracy.
@@ 5 Analysis
Table 3: Memory footprint comparison (GB) for Llama2 [llama2_2023] models under different quantization methods. LittleBit is evaluated at various BPWs, achieved by adjusting the latent rank $r$. Compression factors relative to FP16 are shown in parentheses.The gap between the effective BPW and the compression factor arises because quantization is applied only to the Transformer blocks, while components like the embedding layer and the lm_head remain in FP16.
Model FP16 BiLLM [billm2024] OneBit [onebit2024] LittleBit (Ours)

### v4

@@ 1 Introduction
Large Language Models (LLMs) based on the Transformer architecture [52] have revolutionized natural language processing. However, their immense scale, reaching hundreds of billions or trillions of parameters  [37, 22], leads to prohibitive computational and memory costs, particularly GPU VRAM, hindering their widespread deployment on consumer or edge devices [66, 38, 63].
Model quantization, which reduces numerical precision, is a key technique to address these bottlenecks [20]. While post-training quantization (PTQ) methods like GPTQ [11] and AWQ [30] compress to $\sim$4 bits, further compression (e.g., to 1-bit) typically requires quantization-aware training (QAT) [23, 27] to maintain performance. Indeed, QAT has demonstrated the ability to enable substantial 1-bit compression while preserving model fidelity [53, 57, 32].
However, even 1-bit models (e.g., $\sim$15.4GB for 70B parameters [20]) can exceed the capacity of highly resource-constrained devices [67], motivating the exploration of sub-1-bit quantization. Although prior work [10] has ventured into this area, maintaining performance at extremely low effective bits (e.g., 0.55 bits per weight (BPW)) while ensuring computational efficiency without restrictive hardware dependencies remains a critical challenge [10]. Consequently, it is crucial to pioneer new techniques that can achieve such profound model compression, aiming for bit rates around 0.1 BPW, while maintaining the integrity of model performance.
@@
Our approach stems from two observations. First, LLM weight matrices often exhibit significant low-rank structure [24, 54]. This suggests that factorization methods like Singular Value Decomposition (SVD) [14] could offer a more stable compression pathway than pruning [17], especially under extreme compression regimes. For instance, at compression ratios exceeding 50%, SVD-based approaches [55] consistently outperform pruning methods like Wanda [48] (see Section˜A.5 for details). Second, binarization inherently causes information loss [15]. Recent high-performing 1-bit methods [57, 25, 29] demonstrate that sophisticated scaling over multiple dimensions (e.g., row and column), is crucial for mitigating this loss and stabilizing performance [29].
@@
Extensive experiments show LittleBit’s superiority over STBLLM [10], the leading sub-1-bit technique, on LLMs from 1.3B to 32B parameters. Notably, LittleBit achieves competitive performance on Llama2-13B at an unprecedented 0.1 BPW (Fig.˜1), surpassing STBLLM at 0.55 BPW. Furthermore, the 32B LittleBit model at just 0.3 BPW maintains strong performance, significantly outperforming STBLLM and setting a new state-of-the-art for sub-1-bit quantization.
@@
  • We provide extensive validation showing that LittleBit significantly outperforms STBLLM [10], the prior leading sub-1-bit quantization technique, on various LLM scales.
@@ 2.1 Binarization and Quantization
Network binarization targets extreme compression and acceleration by constraining weights or activations to $\pm 1$, enabling efficient bitwise operations [21, 12]. Although early works like BinaryConnect [7] and BNNs [21] showed feasibility, they faced significant accuracy degradation [39]. This degradation was partially mitigated by introducing scaling factors [41] and employing the Straight-Through Estimator (STE) [3] to handle the non-differentiable quantization function during training [44].
However, directly applying these early techniques to sensitive large language models (LLMs) typically results in unacceptable performance loss [45]. To address this challenge, LLM-specific quantization strategies have been developed, mainly categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [13]. PTQ adapts pre-trained models, achieving competitive results around 4-bit precision with methods such as GPTQ [11] and AWQ [30], but generally struggles to maintain performance below 2-bit precision [45, 20]. The sub-1-bit regime is particularly challenging for PTQ methods due to severe information loss [9]. Recent work such as STBLLM [10] shows that incorporating structured binarized compression can push PTQ into the sub-1-bit regime, yet a non-negligible accuracy gap remains.
In contrast, QAT integrates the quantization process directly into the training loop. This enables the model to learn and adapt to the low-precision format during training, generally yielding superior performance compared to PTQ at very low bit-widths [31]. Successful QAT approaches for low-bit scenarios often employ sophisticated scaling mechanisms alongside the adaptive training itself, as evidenced by several recent methods [57, 25]. However, despite these advancements, consistently achieving high model fidelity when quantizing to less than 1.0 BPW remains a considerable challenge. This persisting difficulty underscores the need for novel architectures and training strategies specifically designed for such extreme compression regimes.
@@ 2.2 Low-Rank Approximation in Quantization
Beyond quantization, the inherent parameter redundancy within deep neural networks [9], particularly evident in LLMs [34, 18], can be effectively exploited using low-rank approximation methods. Techniques such as Singular Value Decomposition (SVD) [14] allow large weight matrices to be represented as products of smaller factors, offering an alternative compression strategy that can be more resilient than pruning, especially under high compression regimes [17]. Recognizing their potential for complementary benefits, researchers have explored combining low-rank factorization to reduce parameter count and quantization to lower parameter precision [42]. Initial studies applied SVD in various roles within the LLM context: some used activation statistics to guide decomposition [55], others integrated SVD principles into initialization or parameter-efficient updates [61, 62, 57], applied transformations prior to factorization [58], or even employed low-rank structures to model quantization errors [60].
@@
A more integrated approach involves incorporating low-rank concepts directly within the QAT methods, as demonstrated by DL-QAT [26], which jointly learns a representation amenable to both low-rank approximation and low-precision quantization, albeit limited to 3-bit quantization. This direction of optimizing structural properties such as rank and numerical precision during training represents a promising path towards more effective LLM compression, with potential for further exploration of even lower precision quantization.
@@ 3.1 LittleBit Architecture
The LittleBit architecture, depicted in Fig.˜2, redesigns linear layers for extreme compression (e.g., 0.1 BPW) by synergistically combining low-rank factorization and binarization. LittleBit first leverages the observed low-rank structure in LLM weights [56] to approximate the weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ as:
@@
To compensate for the significant information loss incurred by binarization [15], LittleBit incorporates learnable FP16 scales. Beyond standard row ($\mathbf{h}$) and column ($\mathbf{g}$) scaling [45, 57, 29], it introduces an additional latent scale ($\bm{\ell}\in\mathbb{R}^{r}$). This latent scale is designed to address the factorization structure by learning the relative importance of each of the $r$ latent dimensions, which correspond to the columns of the binarized factors $\mathbf{U}_{\mathrm{sign}}$ and $\mathbf{V}_{\mathrm{sign}}$:
@@ 4.1 Settings
We evaluate LittleBit on diverse LLM families including Llama [50], Llama2 [51], Llama3 [16], OPT [64], Phi-4 [1], and QwQ [49], spanning sizes from 1.3B to 32B parameters. The primary performance metric is perplexity (PPL), evaluated on the WikiText-2 [35] validation dataset. Additional results on the C4 [40] and PTB [33] datasets are provided in Section˜A.3. We also assess zero-shot accuracy on common reasoning benchmarks: WinoGrande [43], OpenBookQA (OBQA) [36], HellaSwag [59], BoolQ [5], ARC-Easy (ARC-e), ARC-Challenge (ARC-c) [6], and PIQA [4].
@@
The LittleBit model parameters, initialized using the Dual-SVID method (Section˜3.2), are optimized using QAT with knowledge distillation (KD) [19, 31, 65]. The original pre-trained full-precision model serves as the teacher ($\mathcal{T}$) for the LittleBit student model ($\mathcal{S}$). The QAT objective combines the standard output Kullback-Leibler (KL) divergence loss $\mathcal{L}_{\mathrm{out}}$ and an intermediate layer Mean Squared Error (MSE) loss $\mathcal{L}_{\mathrm{inter}}$ to match hidden representations, weighted by an empirically determined $\lambda=10$:
@@
Following the setup in [25], the training data combine WikiText-2 and selected partitions from C4. We use 2048 token sequence length, 5 epochs, Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.999$[28], and cosine LR decay with 2% warm-up (see details in  LABEL:appendix:hyper_parameters). For models employing Grouped-Query Attention (GQA) [2] (e.g., Llama3, Phi-4, QwQ), the latent ranks $r$ for key(K) and value (V) projection layers were specifically adjusted (see Section˜A.2.3) to maintain performance under extreme quantization. For backpropagation, SmoothSign (forward: $\mathrm{sign}(x)$; backward gradient: $\tanh(kx)$, $k=100$) is used, which generally performed better than Straight-Through Estimator (STE) [3] (see Section˜A.2.2). While the concept of a proxy gradient is not new, our contribution lies in the specific application and empirical validation of using the derivative of $\tanh(100x)$ as a stable gradient for QAT in the ultra-low-bit regime, which our ablation study confirms is effective (Section˜A.2.2).
@@
We benchmark LittleBit across a range of effective bits from approximately 1.0 down to 0.1 bits per weight (BPW). In the challenging sub-1-bit regime, we compare against STBLLM [10], a post-training quantization (PTQ) method employing N:M sparsity. For the $\sim$1.0 BPW setting, we compare against the 1-bit PTQ method BiLLM [20] and quantization-aware training (QAT) approaches (OneBit [57], BinaryMoS [25]). Results from standard low-bit methods like GPTQ [11] and OmniQuant [46] are included for broader context.
Table 1: Perplexity (PPL) comparison on WikiText-2 across various LLMs and quantization methods. Lower PPL indicates better performance. BPW denotes effective bits per weight. Methods marked with use quantization-aware training (QAT). STBLLM [10] utilizes N:M sparsity (ratio in parentheses). LittleBit demonstrates strong performance, particularly in the sub-1-bit regime.
@@
Table 2: Zero-shot accuracy (%) comparison on common sense reasoning benchmarks. Compares Full Precision (FP16) models against STBLLM [10] and LittleBit at 0.55 and 0.3 BPW.
@@ 4.2 Main Results
Table˜1 shows LittleBit’s perplexity (PPL) results against baselines. LittleBit consistently performs robustly, especially in the sub-1-bit regime. Compared to STBLLM [10], LittleBit achieves markedly superior PPL at 0.8, 0.7, and 0.55 BPW. For instance, on Llama2-7B, LittleBit at 0.55 BPW achieves a PPL of 10.47, a substantial improvement over STBLLM’s 30.67. Similar significant gains are observed across other models and BPWs; for Llama2-13B at 0.8 BPW, LittleBit records a PPL of 8.52 compared to STBLLM’s 11.90.
@@
A key strength of LittleBit is its exceptional stability and performance in the extreme sub-0.5 BPW range, specifically at 0.3 BPW and 0.1 BPW. In this territory, methods like STBLLM exhibit severe performance degradation (e.g., Llama2-7B PPL of 1.8e3 at 0.3 BPW). In contrast, LittleBit maintains strong PPL scores, such as 12.00 for Llama2-7B at 0.3 BPW and 15.92 at an unprecedented 0.1 BPW. When configured for an effective bits around 1.0 BPW by adjusting the latent rank $r$, LittleBit delivers perplexity comparable to specialized 1-bit QAT methods like OneBit [57] (e.g., Llama2-7B: LittleBit 9.08 vs. OneBit 8.36) and PTQ methods such as BiLLM [20]. While a method like BinaryMoS [25] may exhibit slightly lower PPL in some 1-bit scenarios (e.g., Llama2-7B: BinaryMoS 7.74), this can be attributed to its use of more complex mechanisms like dynamic scaling, which represent a different set of architectural and computational trade-offs. Analyzing the trade-off by adjusting latent rank to tune effective bits from 1.0 down to 0.1 BPW reveals that while performance degradation is minimal down to  0.55 BPW, a quantization cliff appears between 0.3-0.1 BPW, positioning the 0.3-0.55 BPW range as an optimal sweet spot for balancing compression and accuracy.
@@ 5 Analysis
Table 3: Memory footprint comparison (GB) for Llama2 [51] models under different quantization methods. LittleBit is evaluated at various BPWs, achieved by adjusting the latent rank $r$. Compression factors relative to FP16 are shown in parentheses.The gap between the effective BPW and the compression factor arises because quantization is applied only to the Transformer blocks, while components like the embedding layer and the lm_head remain in FP16.
Model FP16 BiLLM [20] OneBit [57] LittleBit (Ours)